OpenStudy (sleepyjess):

Simplify the expression. \(\dfrac{(csc~y~+~cot~y)(csc~y~-~cot~y)}{csc~y}\) -sin y cos y sin y \(\leftarrow\) I think this one csc y

myininaya (myininaya):

(a+b)(a-b)=a^2-b^2

myininaya (myininaya):

and i remember that sin^2(x)+cos^2(x)=1 so 1+cot^2(x)=csc^2(x) which implies 1=csc^2(x)-cot^2(x)

OpenStudy (anonymous):

you are right

OpenStudy (sleepyjess):

So it would be \(\dfrac{csc^2y-cot^2y}{cscy}\)?

myininaya (myininaya):

yeah

myininaya (myininaya):

then apply the pythagorean identity i just mentioned

myininaya (myininaya):

are however you spell that one guy's name

OpenStudy (sleepyjess):

then \(\dfrac{1}{cscy}=siny\)

myininaya (myininaya):

yep yep

OpenStudy (sleepyjess):

Thanks! @myininaya and @satellite73